# RoadFormer: Duplex Transformer for RGB-Normal Semantic Road Scene Parsing

### v3

@@ I Introduction
Existing road scene parsing approaches predominantly fall under two categories: geometry-based and data-driven ones [1]. The algorithms in the former category typically leverage explicit geometric models to represent regions of interest (RoIs), and then proceed to optimize specific energy functions for accurate RoI extraction. For instance, the study presented in [7] employs a B-spline model to fit the road disparity map, which is subsequently projected onto a 2D v-disparity histogram for freespace detection. Additionally, the research presented in [8] introduces a disparity map transformation algorithm designed specifically for the effective detection of road defects. Nonetheless, actual roads are often uneven, rendering such approaches occasionally infeasible [ma2022computer].
@@ IV-B Experimental Setup and Evaluation Metrics
Dataset
Backbone
IoU (%) Fsc (%) Pre (%) Rec (%)
SYN-UDTIRI
ConvNeXt
93.38 96.58 96.58 96.74
Swin
93.18 96.47 96.59 96.35
CityScapes
ConvNeXt
95.80 97.86 97.74 97.97
Swin
94.69 97.27 97.25 97.29
@@
SEB
HFFM
FFRM
IoU (%) Acc (%) Fsc (%) Pre (%) Rec (%) FPS
$\times$
$\times$
$\times$
95.11 96.87 97.50 98.12 96.87 21.80
$\checkmark$
$\times$
$\times$
95.45 97.40 97.67 97.95 97.40 21.60
$\times$
$\times$
$\checkmark$
95.49 97.59 97.69 97.79 97.59 21.60
$\times$
$\checkmark$
$\times$
95.34 97.67 97.61 97.55 97.67 20.50
$\times$
$\checkmark$
$\checkmark$
95.80 97.97 97.86 97.74 97.97 20.10
@@ IV-D Experimental Results
Method
IoU (%)
Fsc (%)
Pre (%)
Rec (%)
mIoU (%)
RGB
Mask2Former
93.84
96.82
97.14
96.51
74.80
SegFormer
93.98
96.90
96.02
97.79
64.51
DeepLabv3+
93.82
96.81
96.99
96.63
68.66
HRNet
94.06
96.94
96.29
97.59
70.10
RGB-Normal
FuseNet
91.60
95.60
96.00
95.30
52.70
SNE-RoadSeg
93.80
96.80
96.10
97.50
53.40
RTFNet
94.10
96.90
96.30
97.60
49.60
OFF-Net
89.60
94.50
93.40
95.70
39.20
MFNet
92.10
95.90
94.10
97.70
49.30

### v4

@@ I Introduction
Existing road scene parsing approaches predominantly fall under two categories: geometry-based and data-driven ones [1]. The algorithms in the former category typically leverage explicit geometric models to represent regions of interest (RoIs), and then proceed to optimize specific energy functions for accurate RoI extraction. For instance, the study presented in [7] employs a B-spline model to fit the road disparity map, which is subsequently projected onto a 2D v-disparity histogram for freespace detection. Additionally, the research presented in [8] introduces a disparity map transformation algorithm designed specifically for the effective detection of road defects. Nonetheless, actual roads are often uneven, rendering such approaches occasionally infeasible [9].
@@ IV-B Experimental Setup and Evaluation Metrics
Dataset Backbone IoU (%) Fsc (%) Pre (%) Rec (%)
SYN-UDTIRI ConvNeXt 93.38 96.58 96.58 96.74
Swin 93.18 96.47 96.59 96.35
CityScapes ConvNeXt 95.80 97.86 97.74 97.97
Swin 94.69 97.27 97.25 97.29
@@
SEB HFFM FFRM IoU (%) Acc (%) Fsc (%) Pre (%) Rec (%) FPS
$\times$ $\times$ $\times$ 95.11 96.87 97.50 98.12 96.87 21.80
$\checkmark$ $\times$ $\times$ 95.45 97.40 97.67 97.95 97.40 21.60
$\times$ $\times$ $\checkmark$ 95.49 97.59 97.69 97.79 97.59 21.60
$\times$ $\checkmark$ $\times$ 95.34 97.67 97.61 97.55 97.67 20.50
$\times$ $\checkmark$ $\checkmark$ 95.80 97.97 97.86 97.74 97.97 20.10
@@ IV-D Experimental Results
Method IoU (%) Fsc (%) Pre (%) Rec (%) mIoU (%)
RGB Mask2Former 93.84 96.82 97.14 96.51 74.80
SegFormer 93.98 96.90 96.02 97.79 64.51
DeepLabv3+ 93.82 96.81 96.99 96.63 68.66
HRNet 94.06 96.94 96.29 97.59 70.10
RGB-Normal FuseNet 91.60 95.60 96.00 95.30 52.70
SNE-RoadSeg 93.80 96.80 96.10 97.50 53.40
RTFNet 94.10 96.90 96.30 97.60 49.60
OFF-Net 89.60 94.50 93.40 95.70 39.20
MFNet 92.10 95.90 94.10 97.70 49.30